OpenStudy (anonymous):

log{1/3(a+b)} = 1/2 {log(a) + log(b)} find the value of a^2 + b^2 help me

OpenStudy (anonymous):

a) 9ab b) 7ab c) 5ab d) none of these

myininaya (myininaya):

\[\log(\frac{1}{3}(a+b))=\frac{1}{2}(\log(a)+\log(b))\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can't wait...

myininaya (myininaya):

i don't see how to do it right a way give me a sec to think or minute

OpenStudy (anonymous):

ok no problem

OpenStudy (anonymous):

i got 4ab

OpenStudy (anonymous):

but i am an idiot wrong

OpenStudy (anonymous):

7ab is right

OpenStudy (anonymous):

ready?

myininaya (myininaya):

yes i got 7ab too

OpenStudy (anonymous):

but how

OpenStudy (anonymous):

\[\log(\frac{1}{3}(a+b))=\log(\sqrt{a})+\log(\sqrt{b})\]

myininaya (myininaya):

its a little sloppy

OpenStudy (anonymous):

hmn,

OpenStudy (anonymous):

i don't think that u r correct satellite , why u wrote sqrt_{a} and sqrt_{b}

OpenStudy (anonymous):

who solved this myinianya ?

myininaya (myininaya):

i solved it in that pdf

OpenStudy (anonymous):

myinianya do u have scanner

myininaya (myininaya):

yes

OpenStudy (anonymous):

how ? , is there any software

OpenStudy (anonymous):

exponentiate both sides get \[\frac{1}{3}(a+b)=e^{\log(\sqrt{a})+{\log(\sqrt{b})}}\] \[\frac{1}{3}(a+b)=e^{\log(\sqrt{a})}\times e^{{\log(\sqrt{b})}}\]

myininaya (myininaya):

doesn't everyone lol i don't my husband does all the computer stuff that makes me happy

myininaya (myininaya):

interesting satellite i didn't go that route

OpenStudy (anonymous):

ok , u r my aunt . i am only 13 years old

OpenStudy (anonymous):

so \[\frac{1}{3}(a+b)=\sqrt{a}\sqrt{b}\] or \[(a+b)=3\sqrt{a}\sqrt{b}\]

myininaya (myininaya):

you aunt ?

OpenStudy (anonymous):

square both sides get \[a^2+b^2+2ab=9ab\]

OpenStudy (anonymous):

so \[a^2+b^2=7ab\]

OpenStudy (anonymous):

THERE DID IT!

OpenStudy (anonymous):

no , i meant that u r my aunt .

OpenStudy (anonymous):

thanks satellite , u r much different

myininaya (myininaya):

how do you know me? how are we related?

myininaya (myininaya):

i'm confused is satellite your uncle?

OpenStudy (anonymous):

sorry it took so long to type in

OpenStudy (anonymous):

I don't know u . i meant that u r of age of my aunt :)

OpenStudy (anonymous):

yeah kush is the son of a sister my mother never had

OpenStudy (anonymous):

satellite u r gr8 . thanks very much

myininaya (myininaya):

im 26 you make me sound old by calling me your aunt

myininaya (myininaya):

am i old?

OpenStudy (anonymous):

oh sorry ,then u r didi ( sister's age )

myininaya (myininaya):

lol

OpenStudy (anonymous):

r u on faceboob . myininaya please please

OpenStudy (anonymous):

facebook * add me as friend there please https://www.facebook.com/kushashwaravi.shrimali